A room has three lightbulbs. Each one has a 10% probability of burning out within the month. Write the probability as a percentage rounded to one decimal place.
What is the probability that all three will burn out within the month?

Found 2 solutions by ikleyn, rothauserc:
Answer by ikleyn(52803)

.
0.1*0.1*0.1 = 0.001.


Answer by rothauserc(4718)

10% probability = 0.1
:
Probability that all three fail in a month = (0.1)^3 = 0.001
: